MICROWAVE SPECTROSCOPIC INVESTIGATION OF HNO_{3}⋯(H_{2}O)_{2}

\underline{M. B. CRADDOCK}, C. S. BRAUER, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455; M. C. ORILALL, Department of Chemistry, Concordia University, Moorhead, MN 56562.

Nitric acid is an important reactive species in the atmosphere and the study of its hydrates is of considerable interest. We report the observation of the 1:2 complex HNO_{3}⋯(H_{2}O)_{2} via Fourier transform microwave spectroscopy. A-type spectra for a total of 18 isotpomers were recorded, including ^{15}N, and several H_{2}^{18}O and deuterium containing species. No b-type transitions were found despite calculations predicting a significant dipole moment along the b-principal axis. Spectral splittings observed indicate internal motion of one or both water units within this complex. The HNO_{3}⋯(H_{2}O)_{2} system adopts a cyclic structure in which the second water unit inserts into the weak, secondary hydrogen bond of HNO_{3}⋯H_{2}O, previously studied in our laboratory. The near-linear hydrogen bond between the acidic proton and the closest water unit is 1.632 (16) Å, a contraction of 0.15 Å relative to HNO_{3}⋯H_{2}O. The O⋯O distance between the hydroxyl unit of the acid and the closest water unit is 2.625(16)Å. Detailed structural analysis, discussion of internal dynamics, and comparison to ab initio calculations^{a} will be presented. Structural characterization of HNO_{3}⋯(H_{2}O)_{2} will be discussed in the context of proton transfer, as complexes of this nature help answer the fundamental question of how many water molecules are required to ionize a simple mineral acid.

aP. R. McCurdy, W. P. Hess, S. S. Xantheas J. Phys. Chem. A \underline{106}(33), 7628, (2002).